# Effect of spin-dependent tunneling in a MoSe2/Cr2Ge2Te6 van der Waals heterostructure on exciton and trion emission

### v1

@@ III.2.2 Polar magneto-optical Kerr effect
The polar MOKE measurements were performed in the setup described above using the 1.96 eV diode laser. Its intensity was modulated using a flywheel chopper. A beamsplitter cube was introduced into the detection beam path, so that the reflected laser light could be guided to an optical bridge detector with a pair of balanced photodiodes (see [32] for details). The difference signal from the photodiodes was detected using a lock-in amplifier. MOKE loops were measured by tracking the difference signal as a function of the applied magnetic field controlled by the bipolar current source.
@@ IV Acknowledgements
The authors gratefully acknowledge technical assistance by E. Moldt and fruitful discussions with R. Edhib. S. D. acknowledges financial support by the Humboldt foundation and a startup funding grant provided by the DFG via SPP2244. The authors also acknowledge financial support by the DFG via the following grants: SFB1477 (project No. 441234705), SPP2244 (project No. 443361515) and KO3612/7-1 (project No. 467549803). K. W. and T. T. acknowledge support from the JSPS KAKENHI (grant numbers 20H00354 and 23H02052).

### v2

@@ III.2.2 Polar magneto-optical Kerr effect
The polar MOKE measurements were performed in the setup described above using the 1.96 eV diode laser applying an excitation density of about 1 kW/cm2. All measurements shown in the manuscript were performed at a nominal temperature of about 5 K. Laser intensity was modulated using a flywheel chopper. A beamsplitter cube was introduced into the detection beam path, so that the reflected laser light could be guided to an optical bridge detector with a pair of balanced photodiodes (see [50] for details). The difference signal from the photodiodes was detected using a lock-in amplifier. MOKE loops were measured by tracking the difference signal as a function of the applied magnetic field controlled by the bipolar current source.
@@ IV Acknowledgements
The authors gratefully acknowledge technical assistance by E. Moldt and fruitful discussions with R. Edhib. S. D. acknowledges financial support by the Humboldt foundation and a startup funding grant provided by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via SPP2244. T.K. acknowledges financial support by the DFG via the following grants: SFB1477 (project No. 441234705), SPP2244 (project No. 443361515), KO3612/7-1 (project No. 467549803) and KO3612/8-1 (project No. 549364913). K. Z. and J. F. acknowledge funding by the DFG via SFB 1277 (Project No. 314695032) and SPP 2244 (Project No. 443416183), as well as the European Union Horizon 2020 Research and Innovation Program under contract number 881603 (Graphene Flagship) and FLAGERA project 2DSOTECH. K.W. and T.T. acknowledge support from the JSPS KAKENHI (grant numbers 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan.